OpenStudy (anonymous):

Regan is trying to find the equation of a quadratic that has a focus of (−2, 5) and a directrix of y = 13. Describe to Regan your preferred method for deriving the equation. Make sure you use Regan's situation as a model to help her understand.

OpenStudy (anonymous):

I don't quite understand how to find a when writing the equation.. that's all I need help with.

OpenStudy (anonymous):

@jchick

OpenStudy (jchick):

Sorry it took so long to reply lol

OpenStudy (anonymous):

its cool XD

OpenStudy (anonymous):

I waited till I saw you weren't actively helping someone

OpenStudy (jchick):

The definition of a parabola is that the distance of any point on it from the focus is equal to the distance of the point from the directrix. So if (x,y) is a point on the parabola, then find the distance of (x,y) from the focus (-2,5). Find the distance of (x,y) from y = 13. Equate them and you will have your quadratic equation.

OpenStudy (jchick):

(x+2)^2+(y−5)^2=(y−13)^2 Simplify above.

OpenStudy (anonymous):

but how do you find (x,y)?

OpenStudy (jchick):

no matter what value x is, y will always be 13

OpenStudy (anonymous):

why?

OpenStudy (jchick):

directrix y = 13 can be represented as the point (x, 13)

OpenStudy (anonymous):

oh okay, that makes sense.

OpenStudy (jchick):

So after inputting the points, you have: (x−(−2))2+(y−5)2−−−−−−−−−−−−−−−−−−√=(x−x)2+(y−13)2−−−−−−−−−−−−−−−−√

OpenStudy (jchick):

You can then square both sides to get: (x−(−2))2+(y−5)2= (x−x)2+(y−13)2

OpenStudy (jchick):

Does this make sense so far?

OpenStudy (anonymous):

im not sure... itdoes a little bit

OpenStudy (jchick):

Ok where do you need explanation?

OpenStudy (jchick):

You can then square both sides to get: (x−(−2))^2+(y−5)^2= (x−x)^2+(y−13)^2

OpenStudy (jchick):

sorry that is the correct version

OpenStudy (anonymous):

what are you plugging them into?

OpenStudy (jchick):

What do you mean?

OpenStudy (jchick):

Where?

OpenStudy (anonymous):

like are you plugging those into an equation??

OpenStudy (jchick):

And then that simplifies to (x+2)^2+(y−5)^2=(y−13)^2

OpenStudy (jchick):

Yes

OpenStudy (anonymous):

okay, what equation?

OpenStudy (jchick):

(x−(−2))2+(y−5)2−−−−−−−−−−−−−−−−−−√= (x−x)2+(y−13)2−−−−−−−−−−−−−−−−√

OpenStudy (anonymous):

im so confused....

OpenStudy (jchick):

inserted points (-2,5) and (x, 13) the distance formula: (x−x1)2+(y−y2)2−−−−−−−−−−−−−−−−−√=(x−x2)2+(y−y2)2−−−−−−−−−−−−−−−−−√

OpenStudy (jchick):

Do you understand now?

OpenStudy (anonymous):

like I was learning it like this... The graph of the quadratic is a parabola. The directrix is horizontal, so the parabola is vertical. The focus lies below the directrix, so the parabola opens downwards. General equation of a down-opening parabola:    y = a(x - h)² + k with    a < 0    vertex (h, k)    focal length p = | 1/(4a) |    focus (h, k-p)    directrix y = k+p Plug your data into the equations and solve for a now im really confused... when I was only confused about how to figure out what a is XD

OpenStudy (jchick):

Oh ok

OpenStudy (jchick):

Sorry for confusion

OpenStudy (anonymous):

its cool I just didn't know what you were plugging it all into.. XD

OpenStudy (jchick):

Having internet trouble please hang on a second

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I've never seen this done this way. Pretty cool @jchick It looks like this is using the definition of a parabola as the set of points equidistant between the focus and directrix. A circle is a set of point equidistant from the same point. The equation for a circle is \[(x-h)^2+(y-k)^2=r^2\]. jchick set the equations for 2 circles equal to each other, one for the focus with (-2, 5) and the other for the directrix (x, 13). That's where this came from \[(x+2)^2+(y-5)^2=(x-x)^2+(y-13)\] If you mulitply everything out, and then complete the square you can find a with this method

OpenStudy (anonymous):

so do you use the distance formula to find a?

OpenStudy (anonymous):

that's pretty much what it works out to being. and you don't actually need to complete the square. a will be the coefficient of x²

OpenStudy (jchick):

Sorry I am back and thank you @peachpi

OpenStudy (anonymous):

just FOIL out the equation and you'll have an expanded equation for the parabola

OpenStudy (anonymous):

im still really confused... what is r in that equation??

OpenStudy (anonymous):

Maybe try a different method since this is confusing. We know that the vertex is halfway between the focus and directrix, so it's (-2, 9) and the parabola opens down like you said above. So the focal length is 9 - 5 = 4